A landscaping company needs 200 gallons of 24% fertilizer to fertilize the shrubs in an office park. They have in stock 30% fertilizer and 20% fertilizer. How much of each type should they mix together?
Answer by lwsshak3(11628)

A landscaping company needs 200 gallons of 24% fertilizer to fertilize the shrubs in an office park. They have in stock 30% fertilizer and 20% fertilizer. How much of each type should they mix together?
**
let x=amt of 30% fertilizer to mix
200-x=amt of 20% fertilizer to mix
30%x+20%(200-x)=24%*200
.3x+40-.2x=48
0.x=8
x=80
200-x=120
amt of 30% fertilizer to mix=80 gal
amt of 20% fertilizer to mix=120 gal